# Entropy formula for surface diffeomorphisms

### v1

@@ 2.1 Mechanism behind convex bounds for volume growth
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$. There is some $\alpha\in[0,1]$ such that
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}(f^{n}\sigma)\leq\alpha\cdot h_{\rm top}(f)+(1-\alpha)\cdot\frac{\lambda^{+}(f)}{r}.$
Equivalently,
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}(f^{n}\sigma)\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}.$
Here the supremum is taken over all $C^{r}$ embedded curves $\sigma:[0,1]\to M$ with $\left\|\sigma\right\|_{C^{r}}\leq 1$ and $\left\|d_{t}\sigma\right\|\geq 1/2,\,t\in[0,1]$.
@@
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$. There is some $\alpha\in[0,1]$ such that
@@
Equivalently,
The reason for presenting Theorems 2.1 and 2.2 in a convex-combination form is that our proof actually yields a more refined upper bound, in which the coefficient $\alpha\in[0,1]$ reflects how the volume growth is distributed between the entropic contribution and the Lyapunov-type contribution (the Yomdin term). In particular, the argument implicitly determines an admissible choice of $\alpha$ and thus contains strictly more information than the crude bound given by the maximum. However, making this choice explicit would require introducing a substantial amount of additional technical machinery. For this reason, we do not pursue an explicit description of $\alpha$ here in the statements. As a consequence, the statements of the theorems may be equivalently weakened to the simpler form involving $\max\{h_{\rm top}(f),\lambda^{+}(f)/r\}$.
@@
The key insight behind the convex bound in Theorem 2.1 is that these two sources of complexity need not act at the same time. Instead, the time interval $[0,n)$ can be decomposed into two major types of periods:
@@
During geometric times, one only needs to count the number of dynamical balls that the curve visits, which contributes an entropy-type growth, while no significant local complexity is generated inside each ball. In contrast, during neutral times, the curve stays inside a single dynamical ball, so the global complexity vanishes, and the growth is entirely governed by local effects, captured by the Yomdin term $\lambda^{+}(f)/r$. See Figure 1.
@@
We emphasize that all these works are formulated in the presence of a reference invariant measure, and the quantities under consideration are measure-theoretic invariants, such as metric entropy, Lyapunov exponents, or the uniform length of unstable manifolds w.r.t. some ergodic measures. In contrast, our approach focuses on purely topological quantities, namely the volume growth of sub-manifolds and of the tangent cocycle. The absence of a reference measure leads to several substantial technical difficulties. We mention here two of the main ones.
  • The presence of periodic sources  [^6] A periodic point $p$ is called a periodic source if both Lyapunov exponents of $p$ are positive.. When a reference hyperbolic measure $\mu$ is given, the analysis may be restricted to $\mu$-typical points, which are not influenced by periodic sources. In our setting, however, when considering, for example, the volume growth of a curve $\sigma$, a large portion of points in $\sigma$ may spend a long time near periodic sources. During such intervals, which still correspond to geometric times, the curve needs time to escape from the ”trapping region” of these sources. As a consequence, one must carefully control the amount of topological entropy produced during these geometric times, avoiding an unnecessary decomposition of the curve into exponentially many pieces before it has moved away from periodic sources. To address this issue, we introduce a finer decomposition of time and define the notion of trapping times, which allows us to prevent an excessive creation of entropy (see Proposition 4.7).
  • Lack of uniformity in the absence of a reference measure. Without a reference measure, it becomes more delicate to justify that no local complexity is generated during geometric times in a uniform way. We overcome this difficulty by using, in Proposition 4.9, an upper semi-continuity property for sub-additive sequences, which provides the necessary control.
2.2 Proofs of Theorem A and Theorem C
We now directly apply the two convex bounds established above to prove Theorem A and Theorem C, in order to provide the reader with a clear overview of the global framework.
@@ 2.2 Proofs of Theorem A and Theorem C
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\alpha\cdot h_{\rm top}(f)+(1-\alpha)\cdot\frac{\lambda^{+}(f)}{r}\leq h_{\rm top}(f).$
@@
$\leq\alpha\cdot h_{\rm top}(f)+(1-\alpha)\cdot\frac{\lambda^{+}(f)}{r}$
$\leq h_{\rm top}(f).$
The proof of Theorem C is complete.
Let us remark that Theorem A can also be derived from Theorem C. Indeed, from the proof of Theorem 2.2 given in Section 5.1, one obtains
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}(f^{n}\sigma).$ (7)
More precisely, by a Fubini-type argument (see Section 5.1), there exists a constant $K>0$ such that for any $n\geq 1$,
As a consequence, Theorem C combined with (7) yields
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq h_{\rm top}(f).$
@@ 3.2 Tangent dynamics
$\lim_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-1}\rho\left(\hat{f}^{i}\left(\hat{x}\right)\right)>0.$
@@
  • Let $\mu$ be an invariant measure with $\lambda^{+}(f,x)>0\geq\lambda^{-}(f,x)$ for $\mu$-almost every $x$. Then $\mu$ has a unique unstable lift $\hat{\mu}^{+}$ which can be defined by:
@@
    Moreover (refer to [9, Proposition 3.8],
@@
    For more detailed discussions on the lifted dynamics, we refer to [8, Section 3.4] and the references therein.
@@ 4 Volume growth of the sub-manifolds
  • Proposition 4.7 separates the volume growth into the entropic contribution and the Lyapunov-type contribution (the Yomdin term) via Burguet’s reparametrizations, a variant of Yomdin’s scheme.
@@ 4.1 Reparametrization
The following proposition is a reparametrization result in the spirit of Yomdin theory, developed by Burguet [3, Lemma 12]. It shows that an $\varepsilon$-bounded curve can be decomposed, via affine reparametrizations, into finitely many pieces on which the image under $f$ remains bounded. Moreover, the number of such pieces and the sizes of the reparametrizations are quantitatively controlled by the expansion rates along the curve.
The ceiling function $\left\lceil x\right\rceil$ assigns to each real number $x$ the smallest integer $n$ such that $n\geq x$.
@@
[3, Lemma 12] Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact manifold $M$. There is a small constant $\varepsilon_{f}>0$ such that for any $\varepsilon_{f}$-bounded $C^{r}$ curve $\sigma:[0,1]\to M$ and any two integers $a,b\in\mathbb{Z}$, there is a family $\Theta$ of affine maps from $[0,1]$ to itself with the following properties:
@@
Compared with [3, Lemma 12], we slightly modify the third property by adding the term ” $3r$ ” in the exponent. This modification is acceptable, since the contraction factor is only required to depend on $r$, and any adjustment by a uniform constant depending solely on $r$ does not affect the validity of the result. The role of the additional term ” $3r$ ” is purely technical (see the remarks following Proposition 4.7 for details).
The following result, due to Burguet [4, Lemma 8], can be regarded as an $n$-step and local version of the reparametrization result stated in Proposition 4.1.
For $x\in M$, $n\in\mathbb{N}$ and $\varepsilon>0$, recall the usual $(n,\varepsilon)$-dynamical ball at $x$ is defined by
$B(x,n,\varepsilon):=\{y\in M:\ d\left(f^{k}(y),f^{k}(x)\right)<\varepsilon\ \text{ for all }0\leq k<n\}.$
@@
[4, Lemma 8] Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact manifold $M$. For any integer $p\geq 1$, there are a small constant $\varepsilon_{f^{p}}>0$ and a large constant $B_{p}$ such that for any $\varepsilon_{f^{p}}$-bounded $C^{r}$ curve $\sigma:[0,1]\to M$, any $x\in\sigma$ and any $n\in\mathbb{N}$, there is a family $\Theta$ of affine maps from $[0,1]$ to itself with the following properties:
@@
    $B(x,n,\varepsilon_{f^{p}})\cap\sigma\subset\cup_{\theta\in\Theta}\sigma\circ\theta$,
@@
Compared with [4, Lemma 8], we insert ceiling operations in the third property. This modification is acceptable, since the resulting discrepancy is at most linear in $n$ and can therefore be absorbed into the constant $A_{r}$. Also, to simplify notation, we assume that the constant $A_{r}$ appearing here is the same as the one in Proposition 4.1.
@@ 4.2 Decomposition of the volume growth
$\sum_{i=k}^{n-1}\rho^{\prime}\left(\hat{f}^{k}(\hat{x})\right)\geq n-k.$ (9)
@@
$\sum_{i=k}^{m-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)<m-k.$
Choose $k$ minimal with this property.
  • If $k<n$, since $n$ is a geometric time,
@@
    By the minimality of $k$,
    $\sum_{i=n}^{m-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)\geq m-n.$
    Hence
    $\sum_{i=k}^{m-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)\geq(n-k)+(m-n)=m-k,$
    a contradiction.
  • If $k>n$, since $k$ is not a geometric time, there exists $\ell\in\{0,\dots,k-1\}$ such that
    Adding this to the inequality at $k$ yields with $\ell<k$
    contradicting the minimality of $k$.
Therefore $k=n$, and the conclusion follows. ∎
@@
Let $\varepsilon_{f}\gg\varepsilon_{f^{p}}$ be the small constants in Proposition 4.1 w.r.t. $f,f^{p}$.
A periodic point $p$ of $f$ is called a periodic source if both Lyapunov exponents of $p$ are positive. We list all periodic sources by $z_{1},z_{2},z_{3},\cdots$ (since there are at most countably many of them).
Given a $C^{r}$ curve $\sigma:[0,1]\to M$ and a point $x\in\sigma$, recall that we denote by $\hat{x}=(x,[v])\in\hat{M}$ where $v$ is the unit tangent vector of $\sigma$ at $x$. Given any $m,n\in\mathbb{N}$ with $m\leq n$, any $E\subset[0,n)$ and a family of integers $\beta:=\left\{\left(\beta^{\prime}_{i},\,\beta^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$, we define
$O^{p,m,L}_{n,E,\beta}=\Big\{x\in\sigma:$ $\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,\,\exists 1\leq s\leq n,\text{ s.t. }x\in B(z_{s},m,\frac{\varepsilon_{f}}{100\left\|Df\right\|})\text{ and }$
@@
$O^{p,0,L}_{n,E,\beta}=\Big\{x\in\sigma:$ $\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,\,\forall 1\leq s\leq n,\text{ s.t. }x\notin B(z_{s},\frac{\varepsilon_{f}}{100\left\|Df\right\|})\text{ and }$
@@
  • any $O^{p,m,L}_{n,E,\beta}$ with $\beta:=\{(\beta^{\prime}_{i},\beta^{\prime\prime}_{i})\}_{i\in[0,n)}$.
@@
The set $O^{p,m,L}_{n,E,\beta}$ represents a portion of the curve $\sigma$ on which all points share the same collection of (L-expanded) geometric times $E$ (we call the complementary times the neutral times) and the same derivative data along their orbits. For simplicity, we assume that $\sigma$ intersects only one periodic source, namely a repeller $z$.
@@
It is worth emphasizing that, at neutral times, the curve does not necessarily fail to grow. Rather, by definition of geometric times, the growth rate is modest—specifically, it is bounded by a constant strictly less than $1$ (see the definition of geometric times (9)). The precise threshold is irrelevant for our purposes; what matters is that this bound is uniform and does not depend on the dynamics of $f$. In practice, we absorb this contribution by the additional ”3r3r” term appearing in Proposition 4.1.
@@
Recall that $1\gg\varepsilon_{f}\gg\varepsilon_{f^{p}}>0$ are the small constants in Proposition 4.1 w.r.t. $f$ and $f^{p}$. Let $B_{p}$ be the constant in Proposition 4.3. We may further assume that $\varepsilon_{f^{p}}$ is sufficiently small and $B_{p}$ is sufficiently large such that
@@
    $\log r(f,k,\varepsilon_{f^{p}})\leq B_{p}+k\left(h_{\rm top}(f)+1\right),$
@@
Given $1\leq s\leq n$ and a family of integers $\gamma:=\left\{\left(\gamma^{\prime}_{i},\,\gamma^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$, we define:
$\Omega^{p,m,L,s}_{n,E,\gamma}$$\Big\{$ $x\in\sigma:~\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,x\in B(z_{s},m,\frac{\varepsilon_{f}}{100\left\|Df\right\|})\text{ and }$
@@
    $\log\#\Theta_{i}\leq\#\left([m,a_{i}]\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i}\cdot 2B_{p}+\left(a_{i}+1\right)\cdot A_{r}$
@@
We start an inductive construction to build $\Theta_{i\leq l_{0}}$.
@@
    As a consequence, given any $\theta_{i}\in\Theta_{i}$, after composing with a suitable reparametrization, we may assume  [^14] This follows directly from the definition of $\Omega^{p,m,L,s}_{n,E,\gamma}$, since we only consider points lying in the dynamical ball $B(z_{s},m,\frac{\varepsilon_{f}}{100\left\|Df\right\|})$.
@@
    We apply Proposition 4.1 with $a=\gamma^{\prime}_{a_{i}},b=\gamma^{\prime\prime}_{a_{i}}$, we get a family $\Theta$ of affine maps from $[0,1]$ to itself with
@@
    $f^{a_{i+1}}\circ\sigma\circ\theta_{i}\circ\theta\subset B\left(f^{a_{i+1}}\left(z_{l}\right),\frac{\varepsilon_{f}}{10}\right).$
    Hence we may assume, after composing $\theta$ with a suitable reparametrization  [^15] This can be shown directly from the definitions; see [3, Lemma 13] and the references therein for a precise argument., that $f^{a_{i+1}}\circ\sigma\circ\theta_{i}\circ\theta$ is $\varepsilon_{f}$-bounded. We then define
@@
    $\leq\#\left([m,a_{i+1}]\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i+1}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i+1}\cdot 2B_{p}+\left(a_{i+1}+1\right)\cdot A_{r}.$

### v2

@@ 2.1 Mechanism behind the maximum bounds for volume growth
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$. We have
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}(f^{n}\sigma)\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}$
where the supremum is taken over all $C^{r}$ embedded curves $\sigma:[0,1]\to M$ with $\left\|\sigma\right\|_{C^{r}}\leq 1$ and $\left\|d_{t}\sigma\right\|\geq 1/2,\,t\in[0,1]$.
Remark.
This maximum bound improves the classical estimate in Yomdin theory [35], where only the following sum bound is obtained:
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}(f^{n}\sigma)\leq h_{\rm top}(f)+\frac{\lambda^{+}(f)}{r}.$
@@
Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact surface $M$. We have
The inequalities in Theorems 2.1 and 2.2 can be equivalently expressed in a convex combination form: there is some $\alpha\in[0,1]$ such that
$\limsup_{n\to+\infty}\frac{1}{n}\log\sup_{\sigma}{\rm Vol}(f^{n}\sigma)\leq\alpha\cdot h_{\rm top}(f)+(1-\alpha)\cdot\frac{\lambda^{+}(f)}{r},$ (7)
@@
The reason for presenting this convex-combination form is that our proof actually yields a more refined upper bound, in which the coefficient $\alpha\in[0,1]$ reflects how the volume growth is distributed between the entropic contribution and the Lyapunov-type contribution (the Yomdin term). In particular, the argument implicitly determines an admissible choice of $\alpha$ and thus contains strictly more information than the crude bound given by the maximum. However, making this choice explicit would require introducing a substantial amount of additional technical machinery. For this reason, we do not pursue an explicit description of $\alpha$ here in the statements. As a consequence, the statements of the theorems may be equivalently weakened to the simpler form involving $\max\{h_{\rm top}(f),\lambda^{+}(f)/r\}$.
@@
The key insight behind the convex bound in (7) is that these two sources of complexity need not act at the same time. Instead, the time interval $[0,n)$ can be decomposed into two major types of periods:
@@
During geometric times, one only needs to count the number of dynamical balls that the curve visits, which contributes an entropy-type growth, while no significant local complexity is generated inside each dynamical ball. In contrast, during neutral times, the curve stays inside a single dynamical ball, so the global complexity vanishes, and the growth is entirely governed by local effects, captured by the Yomdin term $\lambda^{+}(f)/r$. See Figure 2.
@@
We emphasize that all these works are formulated in the presence of a reference invariant measure, and the quantities under consideration are measure-theoretic invariants, such as metric entropy, Lyapunov exponents. In contrast, our approach focuses on purely topological quantities, namely the volume growth of the sub-manifolds and of the tangent cocycle. The absence of a reference measure leads to several substantial technical difficulties. We mention here two of the main ones.
  • The presence of periodic sources  [^17] A periodic point $p$ is called a periodic source if both Lyapunov exponents of $p$ are positive.. When a reference hyperbolic measure $\mu$ is given, the analysis may be restricted to $\mu$-typical points, which are not influenced by periodic sources. In our setting, however, when considering, for example, the volume growth of a curve $\sigma$, a large portion of points in $\sigma$ may spend a long time near periodic sources. During such times, which still correspond to geometric times, the curve needs time to escape from the ”trapping region” of these sources. As a consequence, one must carefully control the amount of topological entropy produced during these geometric times, avoiding an unnecessary decomposition of the curve into exponentially many pieces before it has moved away from periodic sources. To address this issue, we introduce a finer decomposition of time and define the notion of trapping times, which allows us to prevent an excessive creation of entropy (see Proposition 4.8).
  • Lack of uniform control on the local complexity at geometric times Without a reference measure, it becomes more delicate to justify that no local complexity is generated during geometric times in a uniform way. We overcome this difficulty by exploiting several partitions of the curve, as introduced in Proposition 4.10, together with an upper semi-continuity property for sub-additive sequences, which yields the required control.
2.2 Proofs of Theorem A and Theorem B
We now directly apply the two bounds established above in Theorem 2.1 and 2.2 to prove Theorem A and Theorem B, in order to provide the reader with a clear overview of the global framework.
@@ 2.2 Proofs of Theorem A and Theorem B
$\limsup_{n\to+\infty}\frac{1}{n}\log\int_{M}\left\|Df^{n}_{x}\right\|\,dx\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}=h_{\rm top}(f).$
@@
$\leq\max\left\{h_{\rm top}(f),\,\,\frac{\lambda^{+}(f)}{r}\right\}$
$=h_{\rm top}(f).$
The proof of Theorem B is complete.
@@ 3.2 Tangent dynamics
$\lim_{n\to+\infty}\frac{1}{n}\sum_{i=0}^{n-1}\rho\left(\hat{f}^{i}\left(\hat{x}\right)\right)>0.$ (9)
@@
  • Given an $\hat{f}-$invariant measure $\hat{\mu}$, if (9) holds for $\hat{\mu}-$almost every point $\hat{x}$, then the Lyapunov exponent $\lambda^{+}(f,x)$ is positive for $\mu$-almost every $x$, where $\mu$ is the projection of $\hat{\mu}$. This follows from the fact that for $\hat{\mu}-$almost every point $\hat{x}=(x,[v])\in\hat{M}$ with $\|v\|=1$,
    $\frac{1}{n}\sum_{i=0}^{n-1}\rho\left(\hat{f}^{i}(\hat{x})\right)=\frac{1}{n}\log\left\|Df_{x}^{n}(v)\right\|.$
@@
  • Let $\mu$ be an $f$-invariant measure with $\lambda^{+}(f,x)>0\geq\lambda^{-}(f,x)$ for $\mu$-almost every $x$. Then $\mu$ has a unique unstable lift $\hat{\mu}^{+}$ which can be defined by:
@@
    Moreover (refer to [10, Proposition 3.8],
@@
    For more detailed discussions on the lifted dynamics, we refer to [9, Section 3.4] and the references therein.
@@ 4 Volume growth of the sub-manifolds
  • Proposition 4.8 separates the volume growth into the entropic contribution and the Lyapunov-type contribution (the Yomdin term) via Burguet’s reparametrizations, a variant of Yomdin’s reparametrizations.
@@ 4.1 Reparametrization
The following proposition is a reparametrization result in the spirit of Yomdin theory, developed by Burguet [4, Lemma 12]. It shows that an $\varepsilon$-bounded curve can be decomposed, via affine reparametrizations, into finitely many pieces on which the image under $f$ remains bounded. Moreover, the number of such pieces and the sizes of the reparametrizations are quantitatively controlled by the expansion rates along the curve.
The ceiling function $\left\lceil x\right\rceil$ assigns to each real number $x$ the smallest integer $n$ such that $n\geq x$. Recall that, given a curve $\sigma$, $\hat{x}:=(x,[v])\in\hat{M}$ where $v$ is the unit tangent vector of $\sigma$ at $x\in\sigma$.
@@
[4, Lemma 12] Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact manifold $M$. There is a small constant $\varepsilon_{f}>0$ such that for any $\varepsilon_{f}$-bounded $C^{r}$ curve $\sigma:[0,1]\to M$ and any two integers $a,b\in\mathbb{Z}$, there is a family $\Theta$ of affine maps from $[0,1]$ to itself with the following properties:
@@
Compared with [4, Lemma 12], we slightly modify the third property by adding the term ” $3r$ ” in the exponent. This modification is acceptable, since the contraction factor is only required to depend on $r$, and any adjustment by a uniform constant depending solely on $r$ does not affect the validity of the result. The role of the additional term ” $3r$ ” is purely technical (see the remarks following Proposition 4.8 for details).
The following result, due to Burguet [5, Lemma 8], can be regarded as an $n$-step and local version of the reparametrization result stated in Proposition 4.1.
For $x\in M$, $n\in\mathbb{N}$ and $\varepsilon>0$, recall the usual $(n,\varepsilon)$ dynamical ball at $x$ is defined by
$B(x,n,\varepsilon):=\left\{y\in M:\ d\left(f^{k}(y),f^{k}(x)\right)<\varepsilon\ \text{ for all }0\leq k<n\right\}.$
@@
[5, Lemma 8] Let $f$ be a $C^{r}$ ($r>1$) diffeomorphism on a compact manifold $M$. For any integer $p\geq 1$, there are a small constant $\varepsilon_{f^{p}}>0$ and a large constant $B_{p}$ such that for any $\varepsilon_{f^{p}}$-bounded $C^{r}$ curve $\sigma:[0,1]\to M$, any $x\in\sigma$ and any $n\in\mathbb{N}$, there is a family $\Theta$ of affine maps from $[0,1]$ to itself with the following properties:
@@
    $B\left(x,n,\varepsilon_{f^{p}}\right)\cap\sigma\subset\cup_{\theta\in\Theta}\sigma\circ\theta$,
@@
Compared with [5, Lemma 8], we insert ceiling operations in the third property. This modification is acceptable, since the resulting discrepancy is at most linear in $n$ and can therefore be absorbed into the constant $A_{r}$. Also, to simplify notation, we assume that the constant $A_{r}$ appearing here is the same as the one in Proposition 4.1.
@@ 4.2 Decomposition of the volume growth
$\sum_{i=k}^{n-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)\geq n-k.$ (11)
@@
$\sum_{i=k}^{m-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)<m-k.$ (12)
Choose $k$ minimal with this property. If $k>n$, since $k$ is not a geometric time, there exists $\ell\in\{0,\dots,k-1\}$ such that
Adding this to inequality (12) yields with $\ell<k$
contradicting the minimality of $k$. Therefore we must have $k\leq n$. If $k=n$, the conclusion holds by inequality (12) (this remains true even in the case $n=0$, where $k$ must also be 0). If $k<n$, since $n$ is a geometric time,
@@
Subtracting this from the inequality (12) yields
$\sum_{i=n}^{m-1}\rho^{\prime}\left(\hat{f}^{i}(\hat{x})\right)<(m-k)-(n-k)=m-n,$
which is exactly the desired conclusion.
@@
Let $\varepsilon_{f}\gg\varepsilon_{f^{p}}$ be the small constants in Proposition 4.1 w.r.t. $f$ and $f^{p}$.
Definition 4.7.
A periodic point $z$ of $f$ is called a periodic source if both Lyapunov exponents of $z$ are positive.
Let $T_{z}$ denote the period of $z$ and let $a:=\min_{i\neq j}\,d\left(f^{i}(z),f^{j}(z)\right)$. We fix a small number $0<r_{z}<\min\{\frac{a}{100},\frac{\varepsilon_{f}}{100\left\|Df\right\|}\}$  [^18] The purpose of introducing the small number $r_{z}$ will be revealed in Proposition 4.10. such that for any $0\leq k\leq T_{z}-1$,
  1. 1.
    $r_{z}=r_{f^{k}\left(z\right)}$,
  2. 2.
    $f^{-T_{z}}\left(B\left(f^{k}(z),r_{z}\right)\right)\subset B\left(f^{k}(z),r_{z}\right),$
  3. 3.
    $f\left(B\left(f^{k}(z),r_{z}\right)\right)\subset B\left(f^{k+1}(z),\frac{a}{100}\right).$
We list all periodic sources by $z_{1},z_{2},z_{3},\cdots$ (since there are at most countably many of them). We may choose $r_{z}$ sufficiently small so that the associated balls $B(z_{i},r_{z_{i}})$ are pairwise disjoint (since periodic sources form a discrete set).
Given a $C^{r}$ curve $\sigma:[0,1]\to M$ and a point $x\in\sigma$, recall that we denote by $\hat{x}=(x,[v])\in\hat{M}$ where $v$ is the unit tangent vector of $\sigma$ at $x$. Given any $L,m,n,p\in\mathbb{N}$ with $m\leq n$, any $E\subset[0,n)$ and a family of integers $\beta:=\left\{\left(\beta^{\prime}_{i},\,\beta^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$, we define
$O^{p,m,L}_{n,E,\beta}=\Big\{x\in\sigma:$ $\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,\,\exists 1\leq s\leq n,\text{ s.t. }x\in B(z_{s},m,r_{z_{s}})\text{ and }$
@@
$O^{p,0,L}_{n,E,\beta}=\Big\{x\in\sigma:$ $\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,\,\forall 1\leq s\leq n,\text{ s.t. }x\notin B(z_{s},r_{z_{s}})\text{ and }$
@@
  • any $O^{p,m,L}_{n,E,\beta}$ with $\beta:=\left\{\left(\beta^{\prime}_{i},\beta^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$.
@@
The set $O^{p,m,L}_{n,E,\beta}$ represents a portion of the curve $\sigma$ on which all points share the same collection of (L-extended) geometric times $E$ (we call the complementary times of $E$ the neutral times) and the same derivative data along their orbits. For simplicity, we assume that $\sigma$ intersects only one periodic source, namely a repeller $z$.
@@
It is worth emphasizing that, at neutral times, the curve does not necessarily fail to grow. Rather, by definition of geometric times, the growth rate is modest-specifically, it is bounded by a constant strictly less than $1$ (see the definition of geometric times (11)). The precise threshold is irrelevant for our purposes; what matters is that this bound is uniform and does not depend on the dynamics of $f$.
@@
Recall that $1\gg\varepsilon_{f}\gg\varepsilon_{f^{p}}>0$ and $A_{r}$ are the constants in Proposition 4.1 w.r.t. $f$ and $f^{p}$. Let $B_{p}$ be the constant in Proposition 4.3. We may further assume that $\varepsilon_{f^{p}}$ is sufficiently small and $B_{p}$ is sufficiently large such that
@@
    $\log r(f,k,\varepsilon_{f^{p}})\leq B_{p}+k\left(h_{\rm top}(f)+A_{r}\right),$
@@
Given $1\leq s\leq n$ and a family of integers $\gamma:=\left\{\left(\gamma^{\prime}_{i},\,\gamma^{\prime\prime}_{i}\right)\right\}_{i\in[0,n)}$, we define  [^25] In the case $m=0$, we define $\Omega^{p,0,L,s}_{n,E,\gamma}$ by replacing the condition $x\in B(z_{s},m,r_{z_{s}})$” in $\Omega^{p,m,L,s}_{n,E,\gamma}$ with $x\notin B(z_{s},r_{z_{s}})$”, as before. :
$\Omega^{p,m,L,s}_{n,E,\gamma}$$\Big\{$ $x\in\sigma:~\left(E(\hat{x})\cap[0,n)\right)^{L}=E,\,x\in B(z_{s},m,r_{z_{s}})\text{ and }$
@@
    $\log\#\Theta_{i}\leq\#\left([m,a_{i}]\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i}\cdot 3B_{p}+a_{i}\cdot 2A_{r}$
@@
We start an inductive construction to build $\Theta_{0\leq i\leq l_{0}}$.
@@
    As a consequence, given any $\theta_{i}\in\Theta_{i}$, after composing with a suitable reparametrization, we may assume  [^27] This follows directly from the definition of $\Omega^{p,m,L,s}_{n,E,\gamma}$, since we only consider points lying in the dynamical ball $B(z_{s},m,r_{z_{s}})$ where $r_{z_{s}}<\frac{\varepsilon_{f}}{100\left\|Df\right\|}$.
@@
    Applying Proposition 4.1 with $a=\gamma^{\prime}_{a_{i}},b=\gamma^{\prime\prime}_{a_{i}}$, we get a family $\Theta$ of affine maps from $[0,1]$ to itself with
@@
    $f^{a_{i+1}}\circ\sigma\circ\theta_{i}\circ\theta\subset B\left(f^{a_{i+1}}\left(z_{s}\right),\frac{\varepsilon_{f}}{10}\right).$
    Hence we may assume, after composing $\theta$ with a suitable reparametrization  [^28] This can be shown directly from the definitions; see [4, Lemma 13] and the references therein for a precise argument., that $f^{a_{i+1}}\circ\sigma\circ\theta_{i}\circ\theta$ is $\varepsilon_{f}$-bounded. We then define
@@
    $\leq\#\left([m,a_{i+1}]\cap E\right)\cdot h_{\rm top}(f)+\sum_{k<a_{i+1}}\frac{\gamma^{\prime}_{k}-\gamma^{\prime\prime}_{k}}{r-1}+\#\Gamma^{2}_{i+1}\cdot 3B_{p}+a_{i+1}\cdot 2A_{r}.$